OpenStudy (anonymous):

Which of the following is equivalent to |4x + 5| = 88? a--4x + 5 = 88 b-4x + 5 = -88 c-4x + 5 = 88 d-4x + 5 = 88 or 4x + 5 = -88

OpenStudy (radar):

Looks like some spaces need to be placed between the choices and the problem? Kind of confusing or those dashes or minus signs?

OpenStudy (radar):

The way these things are solved usually go something like this. 4x+5=88 4x+5=-88 4x=88-5 4x=-88-5 4x=83 4x=-93 x=83/4 x=-93/4 x=83/4, -93/4

OpenStudy (radar):

Kind of like d. without the - sign preceding the 4!!

jhonyy9 (jhonyy9):

d- is right